0 0
Add a comment Improve this question Transcribed image text
Answer #1


continuity equation

A1*v1 = A2*v2

pi*r1^2*v1 = pi*r2^2*v2

r1^2*v1 = r2^2*v2

v2 = (r1/r2)^2*v1 = 5.84*v1


bernoullis pricilple


P1 + 0.5*rho*v1^2 = P2 + 0.5*rho*v2^2 + rho*g*y


(P1-P2) - rho*g*y = 0.5*rho*(v2^2-v1^2)

0.5*rho*(v2^2-v1^2) = (P1-P2) - rho*g*y


0.5*rho*(5.84^2*v1^2-v1^2) = (P1-P2) - rho*g*y


0.5*rho*(5.84^2-1)*v1^2 = (P1-P2) - rho*g*y


0.5*1000*33.1056*v1^2 = (1.7e5-1.24e5)-(1000*9.8*2.5)

#a) v1 = 1.139 m/s


#b) v2 = 5.84*v1 = 6.65176 m/s

#c)


flow rate = A1*v1 = 3.14*2.9*2.9*1e-4*1.139 = 3*10^-3 m^3/s